carol is 5 times as old as her brother she will be 3 times as old as her brother in two years . how old is carol?
Found 2 solutions by rimjhim.tomar, jpaul6@rediffmail.com:
Answer by rimjhim.tomar(21)

let carol's brother be x yrs old.
A.T.Q.
carol's age=5x
now after 2 yrs,
carol's brother is x+2 yrs old.
and carol's age= 5x+2
A.T.Q.
carol will be three times older than her brother.
so,
3(x+2)=5x+2
3x+6=5x+2
5x-3x=6-2
2x=4
x=2
hence carol's age =5x = 5(2) =10 yrs
-Rimjhim

Answer by jpaul6@rediffmail.com(1)